# Foldings in relatively hyperbolic groups

### v1

@@ 2.5 Trivial Segments of Graphs of Groups
In the statement of the combination theorem (Thm. 3.2) the notion of a trivial segment of an $\mathbb{A}$-path plays a crucial role.
@@
Let $\mathbb{A}$ be a graph of groups. A non-degenerate, simple path $s=(e_{1},\ldots,e_{k})$ in $A$ is called a trivial segment of $\mathbb{A}$, if
@@
    $\operatorname{val}(\omega(e_{l}))=2$ for all $l\in\{1,\ldots,k-1\}$, and
@@
Note that $s$ is a (maximal) trivial segment, if and only if $s^{-1}$ is a (maximal) trivial segment. Denote the number of (unordered) pairs $\{s,s^{-1}\}$ of maximal trivial segments by $N(\mathbb{A})$.
@@
A simple argument shows that there is the following upper bound on $N(\mathbb{A})$.
@@
Let $\mathbb{A}$ be a minimal graph of groups. Let $\mathbb{A}_{1},\ldots,\mathbb{A}_{n}$ be the components of $\mathbb{A}$ which arise from deleting all its edges and vertices with trivial edge or vertex group. Then
$N(\mathbb{A})\leq 3b_{1}(A)+2n,$
@@
Let $\bar{A}$ be the graph obtained from $A$ by replacing every maximal trivial segment of $\mathbb{A}$ with a single edge and every $A_{i}$ with a single vertex.
A simple induction shows, that for any graph $\Gamma$ the number of its edges (counting both orientations) is bounded from above by $6b_{1}(\Gamma)+4k$, where $k$ is the number of its vertices of valence at most $2$.
If $A$ consists of a single loop edge then $N(\mathbb{A})\leq 1$ and $b_{1}(A)=1$, hence the claim is true. Otherwise, every vertex of $\bar{A}$ of valence 1 or 2 corresponds to some $\mathbb{A}_{i}$ and since $N(\mathbb{A})$ is at most half the number of edges of $\bar{A}$, it follows:
$N(\mathbb{A})\leq\frac{1}{2}\lvert E\bar{A}\rvert\leq 3b_{1}(\bar{A})+2\lvert\{v\in V\bar{A}\mid\operatorname{val}(v)\leq 2\}\rvert\leq 3b_{1}(A)+2n$
@@ 3.1 $\pi_{1}$-injective $(G,\mathbb{P})$-carrier graphs of groups
Let $G$ be a group, which is torsion-free, hyperbolic relative to =$\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ and has a symmetric finite generating set $X$. Let $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})$ be an $M$-normal $(G,\mathbb{P})$-carrier graph of groups. Let $C\geq 1$.
@@
Let $\nu:=\nu(G,\mathbb{P},X,M)$ be the constant from Lemma & Definition 1.10. Let $C_{1}:=C+4\nu$, $C_{2}:=C_{1}^{2}+3C_{1}$ and $C^{\prime},K\geq 1$ such that any $(K,C_{2},C_{2})$-local-quasigeodesic in $\operatorname{Cay}(G,X\cup\mathcal{P})$ is a $(C^{\prime},C^{\prime})$-quasigeodesic. Such $C^{\prime}$ and $K$ exist by Théorème 1.4 of Chapitre 3 of [16]. Let $\varepsilon_{1}:=3\nu$, $\varepsilon:=\varepsilon(G,\mathbb{P},X,1,K)$ as in Lemma 1.3, and let $\varepsilon_{2}:=2^{K}(K\varepsilon+\varepsilon_{1})$. Let $L^{\prime}:=\varepsilon_{2}K+8\nu$.
@@
Let $g\in\pi_{1}(\mathbb{A},v_{0})$, $t$ some reduced $\mathbb{A}$-path of minimal $X\cup\mathcal{P}$-length with distinguished $\mathcal{P}$-components representing $g$ and $s$ a realization of $t$. As
$\lvert g\rvert_{X\cup\mathcal{P}}^{\mathcal{A}}=\lvert t\rvert_{X\cup\mathcal{P}}^{\mathcal{A}}=\operatorname{len}(s)$
@@
$\lvert\nu_{\mathcal{A}}(g)\rvert_{X\cup\mathcal{P}}=d_{X\cup\mathcal{P}}(s_{-},s_{+}).$
it suffices to show that that $s$ is a $(K,C_{2},C_{2})$-local-quasigeodesic and therefore a $(C^{\prime},C^{\prime})$-quasigeodesic. Thus one needs to show that any subpath of $s$ of length at most $K$ satifies the conditions spelt out in Lemma 1.18 where the constant $C$ in the statement of Lemma 1.18 is $C_{1}$.
@@
Let $\bar{s}$ be a subpath of $s_{j}$ with $\operatorname{len}(\bar{s})\leq K$ If $\bar{s}$ is a subpath of some geodesic subpath of $s$ corresponding to some essential vertex element of $t$ or to some $g_{e}$ for an edge $e$ of $t$, $\bar{s}$ is itself a geodesic and there is nothing to show. Assuming that this is not the case, Lemma 2.12 can be applied twice to obtain a subdivision $\mathcal{A^{\prime}}$ of $\mathcal{A}$ along at most two free edges and a path $s^{\prime}$ with $d_{X}(s^{\prime}_{-},\bar{s}_{-}),d_{X}(s^{\prime}_{+},\bar{s}_{+})\leq\nu$, which is a realization of some $\mathbb{A}^{\prime}$-path $t^{\prime}$.
@@
$\lvert\nu_{\mathcal{A^{\prime}}}(t)\rvert_{X\cup\mathcal{P}}\geq\frac{\lvert t\rvert_{X\cup\mathcal{P}}^{\mathcal{A^{\prime}}}}{C}-C.$
@@
    $\operatorname{len}_{X}(p),\operatorname{len}_{X}(p^{\prime})\geq\varepsilon_{1}=3\nu$, and
@@
Recall that $t$ is of minimal $X\cup\mathcal{P}$-length and has distinguished $\mathcal{P}$-components. Since $\operatorname{len}_{X}(p),\operatorname{len}_{X}(p^{\prime})\geq 3\nu$, $p$ and $p^{\prime}$ must be of the form $q_{1}qq_{2}$ and $q_{1}^{\prime}q^{\prime}q_{2}^{\prime}$, where $q_{1}$, $q_{2}$, $q_{1}^{\prime}$ and $q_{2}^{\prime}$ are $P_{i}$-paths of $X$-length at most $\nu$ and $q$ and $q^{\prime}$ are distinguished $P_{i}$-edges with $\operatorname{len}_{X}(q),\operatorname{len}_{X}(q^{\prime})\geq\nu$. Moreover, $p$ and $p^{\prime}$ can not be part of the same $P_{i}$-component of $s$, hence $s^{\prime}$ is not a $P_{i}$-path.
By applying Lemma 2.21 to $q_{2}s^{\prime}q_{1}^{\prime}$ at most twice, obtain a path $s^{\prime\prime}$, which is a realization of some reduced $\mathbb{A}$-path $t^{\prime}$ with $t^{\prime}_{-}$ and $t^{\prime}_{+}$ peripheral, which is labeled by an element of $P_{i}$, and which fulfills
$\operatorname{len}_{X}^{p}(s^{\prime\prime})\leq\operatorname{len}_{X}^{p}(q_{2}s^{\prime}q_{1}^{\prime})+6\nu.$
@@
$\lvert t^{\prime}\rvert_{X}=\operatorname{len}_{X}^{p}(s^{\prime\prime})\leq\operatorname{len}_{X}^{p}(q_{2}s^{\prime}q_{1}^{\prime})+6\nu\leq\varepsilon_{2}+8\nu\leq L^{\prime}.$
Hence, $t^{\prime}$ fulfills 1., a contradiction.
@@
Now let $O,O^{\prime}\in\mathbb{O}_{\mathcal{A}}$, $o\in O$, $o^{\prime}\in O^{\prime}$ and $g\in\pi_{1}(\mathbb{A},v_{0})$ such that $\nu_{\mathcal{A}}(gog^{-1}o^{\prime})$ is parabolic. By the above, $gog^{-1}o^{\prime}$ must be a parabolic element of some vertex group of $\mathbb{A}$. Since $\mathcal{A}$ is $M$-normal, it follows that $O=O^{\prime}$. Thus $\nu_{\mathcal{A}}(\mathbb{O}_{\mathcal{A}})$ is an induced structure of $(G,\mathbb{P})$ on $\operatorname{Im}\nu_{\mathcal{A}}$. $\square$
@@ 3.2 Combination Theorem
It is now possible to prove the following combination theorem. It states certain concrete (and often checkable) conditions that guarantee that Proposition 3.1 can be applied to show that the map $\nu_{\mathcal{A}}$ is a quasiisometric embedding.
@@
Let $G$ be a group, which is torsion-free, hyperbolic relative to =$\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ and has finite symmetric generating set $X$. Let $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})$ be an $M$-normal $(G,\mathbb{P})$-carrier graph.
There are $C=C(G,\mathbb{P},X,M,N(\mathbb{A}))$ and $L=L(G,\mathbb{P},X,M,N(\mathbb{A}))$ such that the following holds:
@@
      There is some $\mathbb{A}$-path $t$ with $\lvert t\rvert_{X}^{\mathcal{A}}\leq L$, $t_{-}\in VC_{i}$, $t_{+}$ essential, and $A_{t_{+}}\cap\nu_{\mathcal{A}}(t)^{-1}P_{m_{i}}\nu_{\mathcal{A}}(t)\neq 1$, which passes through some trivial segment of $\mathbb{A}$ exactly once.
@@
$C\geq\max\left\{10,\sqrt{\frac{4}{\delta}},\frac{N(\mathbb{A})+1}{\delta}\right\},$
@@
By assumptions 1. and 2., $t^{\prime}$ is of the form
@@
Since $t^{\prime}$ is almost-simple and therefore reduced, $t_{l}$ is a maximal trivial segment of $\mathbb{A}$ for $l\in\{2,\ldots,N-1\}$. It further follows, that $\mathbb{A}^{\prime}$ has at least $N-1$ pairwise distinct maximal trivial segments. Since $N(\mathbb{A}^{\prime})=N(\mathbb{A})$, this implies $N\leq N(\mathbb{A})+1\leq\delta C$.
@@
By assumption 5., $\lvert t\rvert_{X\cup\mathcal{P}}^{\mathcal{A^{\prime}}}>C\lvert t\rvert_{X\cup\mathcal{P}}^{\mathcal{A^{\prime}}}+C^{2}\geq C^{2}$ and hence
@@
Since $N=0$ implies $L_{t}=0$ it follows that $N\geq 1$. Now, as $\delta C\geq N(\mathbb{A})+1\geq N$ and $C\geq 10$, it follows that
@@
      all vertices of $r$ other than $(r_{2})_{+}$ are non-essential.
@@
It is easy to see by considering a shortest $\mathbb{A}$-subpath with the same property, that if $t$ is not as in 1. or 2., then one of the following must be true:
@@
    Since $t$ is reduced and $e_{l_{i}}=e_{l_{i}+1}^{-1}$ it follows that $a_{l_{i}}\notin\omega_{e_{l_{i}}}(A_{e_{l_{i}}})$ for $i\in\{1,2\}$. Thus $v_{l_{1}},v_{l_{2}}\in VA\subset VA^{\prime}$, and $t^{\prime}$ can be assumed to be an $\mathbb{A}$-path rather than an $\mathbb{A}^{\prime}$-path. It follows, that $e_{l_{1}}$, $e_{l_{1}+1}$, $e_{l_{2}}$ and $e_{l_{2}+1}$ are non-peripheral. Hence, $t^{\prime}$ and $(a_{l_{i}})$ are full $\mathbb{A}^{\prime}$-subpaths of $t$ and therefore $\lvert t^{\prime}\rvert_{X}^{\mathcal{A}}\leq\lvert t\rvert_{X}^{\mathcal{A^{\prime}}}\leq L$ and $\lvert a_{l_{i}}\rvert_{X}\leq\lvert t\rvert_{X}^{\mathcal{A^{\prime}}}\leq L$ for $i\in\{1,2\}$.
@@
    if $f_{l}$ is peripheral. It follows that $\lvert s^{\prime}\rvert_{X}^{\mathcal{A^{\prime}}}\leq\lvert s\rvert_{X}^{\mathcal{A^{\prime}}}$.
@@
Let $L\geq 0$ and $\mathcal{A^{\prime}}$ be some subdivision of $\mathcal{A}$ along at most two free edges. Let $t_{i}=(a_{0}^{i},e_{1},a_{1}^{i},\ldots,a_{l-1}^{i},e_{l},1)$ be almost-simple $\mathbb{A}^{\prime}$-paths for $i\in\{1,2\}$, $v_{0}:=\alpha(e_{1})$, $v_{m}:=\omega(e_{m})$ for $m\in\{1,\ldots,l\}$ and $a\in A_{v_{l}}\setminus\omega_{e_{l}}(A_{e_{l}})$. If
@@
    $[t_{2}^{-1}t_{1}]\notin[\omega_{e_{l}}(A_{e_{l}})]:=\{[(a)]\mid a\in\omega_{e_{l}}(A_{e_{l}})\}$,
@@
As $[t_{2}^{-1}t_{1}]\notin[\omega_{e_{l}}(A_{e_{l}})]$, there is some $m\in\{0,\ldots,l-1\}$ such that
@@
Case 2a: The underlying path of $t$ is of the form $rr^{-1}$ for an almost-simple path $r$ where all vertices of $rr^{-1}$ other than $r_{+}$ are non-essential.
@@
a contradiction. Thus, $t$ fulfills (A1(b)).
@@
Let $t=t_{1}st_{2}$. It follows that
@@
Let $G$ be a group, which is torsion-free, hyperbolic relative to =$\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ and has finite symmetric generating set $X$. Let $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})$ be an $M$-prenormal $(G,\mathbb{P})$-carrier graph.
There are $C=C(G,\mathbb{P},X,M,N(\mathbb{A}))$ and $L=L(G,\mathbb{P},X,M,N(\mathbb{A}))$ such that the following holds:
@@
      There is some $\mathbb{A}$-path $t$ with $\lvert t\rvert_{X}^{\mathcal{A}}\leq L$, $t_{-}\in VC_{i}$, $t_{+}$ essential, and $A_{t_{+}}\cap\nu_{\mathcal{A}}(t)^{-1}P_{m_{i}}\nu_{\mathcal{A}}(t)\neq 1$, which passes through some trivial segment of $\mathbb{A}$ exactly once.
@@
The only point worth mentioning is that 1. of Lemma 2.23 is subsumed in (A1) and that 4. of Lemma 2.23 is subsumed in (A2). $\square$

### v2

@@ 2.5 Trivial Segments of Graphs of Groups
In the statement of the combination theorem (Theorem 3.2) the notion of a trivial segment of an $\mathbb{A}$-path plays a crucial role.
@@
Let $\mathbb{A}$ be a finite graph of groups and $v_{0}\in VA$. A non-degenerate, simple or simple closed path $s=(e_{1},\ldots,e_{k})$ in $A$ is called a trivial segment of $\mathbb{A}$ relative $v_{0}$, if
@@
    $\operatorname{val}(\omega(e_{l}))=2$ and $\omega(e_{l})\neq v_{0}$ for all $l\in\{1,\ldots,k-1\}$, and
@@
Note that $s$ is a (maximal) trivial segment, if and only if $s^{-1}$ is a (maximal) trivial segment. Note moreover that any edge $e\in EA$ with trivial edge group is contained in a unique maximal trivial segment. We denote by $N(\mathbb{A},v_{0})$ the number of (unordered) pairs $\{s,s^{-1}\}$ of maximal trivial segments of $\mathbb{A}$ relative $v_{0}$. For a $(G,\mathbb{P})$-carrier graph $\mathcal{A}$ (with base vertex $v_{0}$) we put $N(\mathcal{A}):=N(\mathbb{A},v_{0})$.
@@
The following lemma provides an upper bound on $N(\mathcal{A})$ for a $(G,\mathbb{P})$-carrier graph $\mathcal{A}$.
@@
Let $\mathbb{A}$ be a finite graph of groups and $v_{0}\in VA$ such that any valence $1$ vertex distinct from $v_{0}$ has a non-trivial vertex group. Let $n$ be the number of free factors of $\mathbb{A}$. Then
$N(\mathbb{A})\leq 3b_{1}(A)+2(n+1),$
@@
Let $\bar{A}$ be the graph obtained from $A$ by replacing every maximal trivial segment of $\mathbb{A}$ relative $v_{0}$ with a single edge and every $A_{i}$ with a single vertex.
A simple induction shows, that for any graph $\Gamma$ the number of edges (counting both orientations) is bounded from above by $6b_{1}(\Gamma)+4k$, where $k$ is the number of vertices of valence at most $2$.
Every vertex of $\bar{A}$ of valence at most 2 corresponds to some $\mathbb{A}_{i}$ or to $v_{0}$. Since $N(\mathbb{A})$ is at most half the number of edges of $\bar{A}$, it follows that
$N(\mathbb{A})\leq\frac{1}{2}\lvert E\bar{A}\rvert\leq 3b_{1}(\bar{A})+2\lvert\{v\in V\bar{A}\mid\operatorname{val}(v)\leq 2\}\rvert\leq 3b_{1}(A)+2(n+1).$
@@ 3.1 $\pi_{1}$-injective $(G,\mathbb{P})$-carrier graphs of groups
Let $G$ be a group, which is torsion-free, hyperbolic relative to $\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ and has a symmetric finite generating set $X$. Let $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})$ be an $M$-normal $(G,\mathbb{P})$-carrier graph of groups. Let $C\geq 1$.
@@
Moreover for any tame $\mathbb{A}$-path with realization $s$ the following hold:
  1. (i)
    $s$ is a $(C^{\prime},C^{\prime})$-quasigeodesic.
  2. (ii)
    If $p$ and $p^{\prime}$ are $\mathcal{P}$-components in $s$ such that $\operatorname{len}_{X}(p),\operatorname{len}_{X}(p^{\prime})\geq L^{\prime}$ then $p$ and $p^{\prime}$ are not connected.
@@
Let $\nu:=\nu(G,\mathbb{P},X,M)$ be the constant from Lemma & Definition 1.10. Let $C_{1}:=C+4\nu$, $C_{2}:=C_{1}^{2}+3C_{1}$ and $C^{\prime},K\geq 1$ such that the following hold:
  1. 1.
    Any $(K,C_{2},C_{2})$-local-quasigeodesic in $\operatorname{Cay}(G,X\cup\mathcal{P})$ is a $(C^{\prime},C^{\prime})$-quasigeodesic.
  2. 2.
    $d_{X\cup\mathcal{P}}(\gamma_{-},\gamma_{+})>1$ for any $(C^{\prime},C^{\prime})$-quasigeodesic $\gamma$ of length at leat K2K-2.
Such $C^{\prime}$ and $K$ exist by Théorème 1.4 of Chapitre 3 of [18]. Let $\varepsilon_{1}:=10\nu$, $\varepsilon:=\varepsilon(G,\mathbb{P},X,1,K)$ as in Lemma 1.3, and let $\varepsilon_{2}:=2^{K}(K\varepsilon+\varepsilon_{1})$. Let $L^{\prime}:=\varepsilon_{2}K+14\nu$.
@@
Let $g\in\pi_{1}(\mathbb{A},v_{0})$, $t$ some tame $\mathbb{A}$-path representing $g$ and $s$ a realization of $t$. The $M$-normality guarantees that $t$ has distinguished $\mathcal{P}$-components. As
$\lvert g\rvert_{X\cup\mathcal{P}}^{\mathcal{A}}\leq\lvert t\rvert_{X\cup\mathcal{P}}^{\mathcal{A}}=\operatorname{len}(s)$
@@
$\lvert\nu_{\mathcal{A}}(g)\rvert_{X\cup\mathcal{P}}\leq d_{X\cup\mathcal{P}}(s_{-},s_{+}).$
it suffices to show that $s$ is a $(K,C_{2},C_{2})$-local-quasigeodesic and therefore a $(C^{\prime},C^{\prime})$-quasigeodesic. Thus one needs to show that any subpath of $s$ of length at most $K$ satifies the conditions spelt out in Lemma 1.19 where the constant $C$ in the statement of Lemma 1.19 is $C_{1}$.
@@
Let $\bar{s}$ be a subpath of $s_{j}$ with $\operatorname{len}(\bar{s})\leq K$ . If $\bar{s}$ is a subpath of some geodesic subpath of $s$ corresponding to some essential vertex element of $t$ or to some $g_{e}$ for an edge $e$ of $t$, $\bar{s}$ is itself a geodesic and there is nothing to show. Assuming that this is not the case, Lemma 2.12 can be applied twice to obtain a subdivision $\mathcal{A^{\prime}}$ of $\mathcal{A}$ along at most two free edges and a path $s^{\prime}$ with $d_{X}(s^{\prime}_{-},\bar{s}_{-}),d_{X}(s^{\prime}_{+},\bar{s}_{+})\leq\nu$, which is a realization of some $\mathbb{A}^{\prime}$-path $t^{\prime}$.
@@
$\lvert\nu_{\mathcal{A^{\prime}}}(t^{\prime})\rvert_{X\cup\mathcal{P}}\geq\frac{\lvert t^{\prime}\rvert_{X\cup\mathcal{P}}^{\mathcal{A^{\prime}}}}{C}-C.$
@@
    $\operatorname{len}_{X}(p),\operatorname{len}_{X}(p^{\prime})\geq\varepsilon_{1}=10\nu$, and
@@
Recall that $t$ has distinguished $\mathcal{P}$-components. Since $\operatorname{len}_{X}(p),\operatorname{len}_{X}(p^{\prime})\geq 10\nu$, $p$ and $p^{\prime}$ must be of the form $q_{1}qq_{2}$ and $q_{1}^{\prime}q^{\prime}q_{2}^{\prime}$, where $q_{1}$, $q_{2}$, $q_{1}^{\prime}$ and $q_{2}^{\prime}$ are $P_{i}$-paths of $X$-length at most $5\nu$ and $q$ and $q^{\prime}$ are distinguished $P_{i}$-edges, see also Remark 2.24.
By applying Lemma 2.25 to $q_{2}s^{\prime}q_{1}^{\prime}$ at most twice, obtain a path $s^{\prime\prime}$, which is a realization of some (not necessarily reduced) $\mathbb{A}$-path $t^{\prime}$ with $t^{\prime}_{-}$ and $t^{\prime}_{+}$ peripheral, which is labeled by an element of $P_{i}$, and which fulfills
$\operatorname{len}_{X}^{p}(s^{\prime\prime})\leq\operatorname{len}_{X}^{p}(q_{2}s^{\prime}q_{1}^{\prime})+4\nu.$
@@
$\lvert t^{\prime}\rvert_{X}=\operatorname{len}_{X}^{p}(s^{\prime\prime})\leq\operatorname{len}_{X}^{p}(q_{2}s^{\prime}q_{1}^{\prime})+4\nu\leq 5\nu+\varepsilon+5\nu+4\nu=\varepsilon_{2}+14\nu\leq L^{\prime}.$
It $t^{\prime}$ is not reduced one obtains a contradiction to the tameness of $t$. Thus $t^{\prime}$ is reduced. Hence, $t^{\prime}$ fulfills 1., a contradiction.
@@
Now let $O,O^{\prime}\in\mathbb{O}_{\mathcal{A}}$, $o\in O$, $o^{\prime}\in O^{\prime}$ and $g\in\pi_{1}(\mathbb{A},v_{0})$ such that $\nu_{\mathcal{A}}(gog^{-1}o^{\prime})$ is parabolic. By the above, $gog^{-1}o^{\prime}$ must be a parabolic element of some vertex group of $\mathbb{A}$. Since $\mathcal{A}$ is $M$-normal, it follows that $O=O^{\prime}$. Thus $\nu_{\mathcal{A}}(\mathbb{O}_{\mathcal{A}})$ is an induced structure of $(G,\mathbb{P})$ on $\operatorname{Im}\nu_{\mathcal{A}}$.
This proves that for the choosen constants one of 1.-3. holds. (i) is implicit in the proof and (ii) follows from item 2. in the choice of the constants $C^{\prime}$ and $K$ and the last claim of Lemma 1.19
$\square$
@@ 3.2 Combination Theorem
It is now possible to prove the following combination theorem. It states provides concrete (and often checkable) conditions that guarantee that Proposition 3.1 can be applied to show that the map $\nu_{\mathcal{A}}$ is a quasiisometric embedding.
@@
Let $G$ be a group, which is torsion-free, hyperbolic relative to $\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ and has finite symmetric generating set $X$. Let $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})$ be an $M$-normal $(G,\mathbb{P})$-carrier graph.
There are $C=C(G,\mathbb{P},X,M,N(\mathcal{A}))$ and $L=L(G,\mathbb{P},X,M,N(\mathcal{A}))$ such that the following holds:
$\nu_{\mathcal{A}}\colon(\pi_{1}(\mathbb{A},v_{0}),d_{X\cup\mathcal{P}}^{\mathcal{A}})\to(G,d_{X\cup\mathcal{P}})$ is a $(C,C)$-quasiisometric embedding and $\nu_{\mathcal{A}}(\mathbb{O}_{\mathcal{A}})$ is an induced structure of $(G,\mathbb{P})$ on $\operatorname{Im}\nu_{\mathcal{A}}$ for any peripheral structure $\mathbb{O}_{\mathcal{A}}$ of $\mathcal{A}$, unless one of the following holds:
@@
      There is some $\mathbb{A}$-path $t$ with $\lvert t\rvert_{X}^{\mathcal{A}}\leq L$, $_{-}\in VC_{i}, t+t_{+}$ essential, and $A_{t_{+}}\cap\nu_{\mathcal{A}}(t)^{-1}P_{m_{i}}\nu_{\mathcal{A}}(t)\neq 1$, which passes through some trivial segment of $\mathbb{A}$ exactly once.
@@
$C\geq\max\left\{10,\sqrt{\frac{4}{\delta}},\frac{N(\mathcal{A})+1}{\delta}\right\},$
@@
By assumptions 2. and 3., $t^{\prime}$ is of the form
@@
Since $t^{\prime}$ is almost-simple and therefore reduced, $t_{l}$ is a maximal trivial segment of $\mathbb{A}$ for $l\in\{2,\ldots,N-1\}$. It follows, that $\mathbb{A}^{\prime}$ has at least N1N-1 pairwise distinct maximal trivial segments. Since $N(\mathcal{A}^{\prime})=N(\mathcal{A})$, this implies $N\leq N(\mathcal{A})+1\leq\delta C$.
@@
By assumption 5., $\lvert t\rvert_{X\cup\mathcal{P}}^{\mathcal{A^{\prime}}}>C\lvert\nu_{\mathcal{A^{\prime}}}(t)\rvert_{X\cup\mathcal{P}}^{\mathcal{A^{\prime}}}+C^{2}\geq C^{2}$ and hence
@@
Since $N=0$ implies $L_{t}=0$ it follows that $N\geq 1$. Now, as $\delta C\geq N(\mathcal{A})+1\geq N$ and $C\geq 10$, it follows that
@@
      all vertices of $r$ distinct from $(r_{2})_{+}$ are non-essential.
@@
It is easy to see by considering a shortest $\mathbb{A}$-subpath of $t$ that is also not not as in 1. or 2., that one of the following must be true:
@@
    Note first that in case (b), i.e. if $e_{l_{i}}=e_{l_{i}+1}^{-1}$ for $i\in\{1,2\}$, it follows that $a_{l_{i}}\notin\omega_{e_{l_{i}}}(A_{e_{l_{i}}})$ as $t$ is reduced. This implies that $e_{l_{i}}$ and $e_{l_{i}+1}$ are not peripheral and that $\omega(e_{l_{i}})$ does no come from a subdivision. Thus in both cases $v_{l_{1}},v_{l_{2}}\in VA\subset VA^{\prime}$, and $t^{\prime}$ can be assumed to be an $\mathbb{A}$-path rather than an $\mathbb{A}^{\prime}$-path. It follows, that $e_{l_{1}+1}$ and $e_{l_{2}}$ (and in case (b) $e_{l_{1}}$ and $e_{l_{2}+1}$) are non-peripheral. Hence, $t^{\prime}$ and $(a_{l_{i}})$ are full $\mathbb{A}^{\prime}$-subpaths of $t$ and therefore $\lvert t^{\prime}\rvert_{X}^{\mathcal{A}}\leq\lvert t\rvert_{X}^{\mathcal{A^{\prime}}}\leq L$ and $\lvert a_{l_{i}}\rvert_{X}\leq\lvert t\rvert_{X}^{\mathcal{A^{\prime}}}\leq L$ for $i\in\{1,2\}$.
@@
    if $f_{l}$ is peripheral. In both cases it follows that $\lvert s^{\prime}\rvert_{X}^{\mathcal{A^{\prime}}}\leq\lvert s\rvert_{X}^{\mathcal{A^{\prime}}}$.
@@
Let $L\geq 0$ and $\mathcal{A^{\prime}}$ be some subdivision of $\mathcal{A}$ along at most two free edges. Let $t_{i}=(a_{0}^{i},e_{1},a_{1}^{i},\ldots,a_{l-1}^{i},e_{l},1)$ be almost-simple $\mathbb{A}^{\prime}$-paths for $i\in\{1,2\}$ (with the same underlying path $(e_{1},\ldots,e_{l})$), $v_{0}:=\alpha(e_{1})$, $v_{m}:=\omega(e_{m})$ for $m\in\{1,\ldots,l\}$ and $a\in A_{v_{l}}\setminus\omega_{e_{l}}(A_{e_{l}})$. If
@@
    $t_{2}^{-1}t_{1}]\notin[\omega_{e_{l}}(A_{e_{l}})]:=\{[(a)]\mid a\in\omega_{e_{l}}(A_{e_{l}})\}$,
@@
As $t_{2}^{-1}t_{1}]\notin[\omega_{e_{l}}(A_{e_{l}})]$, there is some $m\in\{0,\ldots,l-1\}$ such that
@@
Case 2a: The underlying path of $t$ is of the form $rr^{-1}$ for an almost-simple path $r$ where all vertices of $rr^{-1}$ distinct from $r_{+}$ are non-essential.
@@
a contradiction. Thus $t_{1}$ does have an edge with trivial edge group and therefore fulfills (A1(b)).
@@
Let $t_{1}$ and $t_{2}$ be such that $t=t_{1}st_{2}$. It follows that
@@
Let $G$ be a group, which is torsion-free, hyperbolic relative to $\mathbb{P}=\{P_{1},\ldots,P_{n}\}$ and has finite symmetric generating set $X$. Let $(\mathcal{A},((\mathcal{C}_{i},c_{i}))_{1\leq i\leq k})$ be an $M$-prenormal $(G,\mathbb{P})$-carrier graph.
There are $C=C(G,\mathbb{P},X,M,N(\mathcal{A}))$ and $L=L(G,\mathbb{P},X,M,N(\mathcal{A}))$ such that the following holds:
@@
      There is some $\mathbb{A}$-path $t$ with $\lvert t\rvert_{X}^{\mathcal{A}}\leq L$, $_{-}\in VC_{i}, t+t_{+}$ essential, and $A_{t_{+}}\cap\nu_{\mathcal{A}}(t)^{-1}P_{m_{i}}\nu_{\mathcal{A}}(t)\neq 1$, which passes through some trivial segment of $\mathbb{A}$ exactly once.
@@
The only point worth mentioning is that 1. of Lemma 2.27 is subsumed in (A2) and that 4. of Lemma 2.27 is subsumed in (A1(a)). $\square$